# Sparse Mean Field Load Balancing in Large Localized Queueing Systems

### 20 Dec 2023 v1

@@ 1.2.4 Dynamical system model
where each $\mathbb{P}(z_{i}(t+1)\mid z_{N_{i}}(t),a_{N_{i}}(t))$ can be computed by the Kolmogorov equation for continuous-time Markov chains, given that the rate of an arrival at a queue is given by the sum of arrival rates assigned by the agents, $\lambda_{t}^{i}=\lambda(t)(1-a_{i}+\sum_{j\in N_{i}}\frac{a_{j}}{|N_{j}|})$, and the package departure rate is fixed to the serving rate $\alpha>0$.
@@
Each agent is associated with a local stage reward $r_{i}(z_{i},a_{i})$ and according global stage reward $r(\mathbf{z},\mathbf{a})=\frac{1}{N}\sum_{i=1}^{N}r_{i}(z_{i},a_{i})$. The reward is given in terms of a penalty for job drops due to each action $a_{i}$. The objective is then to find the localized policy tuple $\theta$ such that the global reward is maximized, starting from initial state distribution $z_{i}(0)\sim\mu_{0}$. em Should we also define the environment state, which is the state of all queues at time $t$, or is it already part of the global state in some way and can be ignored?
1.3 Sparse Graph Mean Field System
@@
In this section, we will consider limits of large graphs and the behaviour in such systems, by using mean field theory from Lacker et al. (2023). In order to obtain a limiting mean field system, we assume a shared policy $\zeta_{i}=\zeta$ for all agents $i$. This assumption is natural, as it often gives state-of-the-art performance in various RL benchmarks Yu et al. (2022); Christianos et al. (2020); Papoudakis et al. (2021) and also allows immediate application to arbitrary sizes of the finite system, as all schedulers can use the same policy. Furthermore, it will allow to scale up the system at any time by simply adding more schedulers and queues, without retraining a policy.
@@
In contrast to typical mean field games Lasry and Lions (2007) and mean field control Bensoussan et al. (2013), we cannot reduce the entire system exactly to a single representative agent and its probability law. This is because in a sparse setting, the neighbourhood state distribution of any agent remains an empirical distribution of finitely many neighbours and hence cannot follow a law of large numbers into a deterministic mean field distribution. Therefore, the neighbourhood and its state distribution remain stochastic and it is necessary to model the probability law of entire neighbourhoods. The modelling of such graphical neighbourhoods is formalized in the following.
1.3.1 Topological structure
@@
To make sense of the limiting topology of our system formally, we introduce technical details, letting finite systems be given by sequences of (potentially random) finite graphs and initial states $(G_{n},\mathbf{z}_{n})$ converging in probability in the local weak sense to some limiting random graph $(G,\mathbf{z})$. Here, we assume that graphs are of bounded degree, i.e. there exists a finite degree $d$ such that all nodes have at most $d$ neighbours. In other words, we define a sequence of systems of increasing size according to a certain topology, which formalizes the scalable architectural choice of a network structure, such as a ring topology or torus.
@@
To define what one understands as convergence in the local weak sense, define first the space of marked rooted graphs $\mathcal{G}_{*}$, the elements of which essentially constitute a snapshot of the entire system at any point in time. Such a marked rooted graph consists of a tuple $(G,\varnothing,\mathbf{z})\in\mathcal{G}_{*}$, where $G=(\mathcal{N},\mathcal{E})$ is a graph, $\varnothing\in\mathcal{N}$ is a particular node of $G$ (the so-called root node), and $z\in\mathcal{Z}^{\mathcal{N}}$ defines states (”marks”) for each node in $G$, i.e. the current queue filling of queues associated to any agent (node). Denote by $B_{k}(G,\varnothing,\mathbf{z})$ the marked rooted subgraph of vertices in the $k$-hop neighbourhood of the root node $\varnothing$. The space $\mathcal{G}_{*}$ is metrized such that sequences $(G_{n},\varnothing_{n},\mathbf{z}_{n})\to(G,\varnothing,\mathbf{z})\in\mathcal{G}_{*}$ whenever for any $k\in\mathbb{N}$, there exists $n^{\prime}$ such that for all $n>n^{\prime}$ there exists a mark-preserving isomorphism $\phi\colon B_{k}(G_{n},\varnothing_{n},\mathbf{z}_{n})\to B_{k}(G,\varnothing,\mathbf{z})$, i.e. with $z_{ni}=z_{\phi(i)}$ for all nodes $i\in G_{n}$ (local convergence).
@@
We will abbreviate elements $(G_{n},\varnothing_{n},\mathbf{z}_{n}),(G,\varnothing,\mathbf{z})\in\mathcal{G}_{*}$ as $(G_{n},\mathbf{z}_{n}),(G,\mathbf{z})$, and their node sets as $G_{n},G$ whenever it is clear from the context. Then, finally, convergence in the local weak sense is formally defined by $\lim_{n\to\infty}\frac{1}{|G_{n}|}\sum_{i\in G_{n}}f(C_{i}(G_{n}))=\mathbb{E}\left[f(G)\right]$ in probability for every continuous and bounded $f\colon\mathcal{G}_{*}\to\mathbb{R}$, where $C_{i}(G_{n})$ denotes the connected component of $i$ in $G_{n}$. In other words, wherever we randomly look in the graph, there will be little difference between the distribution of the random local system state (including its topology), and of the limiting $G$. This holds true e.g. initially if we initialize all queues as empty or i.i.d., and consider various somewhat regular topologies, see Section 2.1. More details also in Lacker et al. (2023).
1.3.2 System model
@@
As discussed in the prequel, consider sequences of possibly random rooted marked graphs (i.e. finite graphs and initial states) $(G_{n},\mathbf{z}_{n})\to(G,\mathbf{z})$ with $|G_{n}|$ agents $i=\{1,\ldots,|G_{n}|\}$, converging in the local weak sense to the potentially infinite-sized system $(G,\mathbf{z})$. For a moment, consider a decentralized, stochastic control policy $\zeta\colon\mathcal{Z}\times\mathcal{T}\to[0,1]$ such that an agent $i\in G$ chooses to offload to a random neighbour (action $1$) with probability $\zeta_{t}(z_{i}(t))$, depending on its own queue state $z_{i}(t)$ only. We can then consider the probability law $\mathcal{L}(G,\mathbf{z})$ of the limiting system $(G,\mathbf{z})$ as the state of a single-agent MDP, similar to the MFC MDP formalism in standard MFC Pham and Wei (2018). At least formally, we do so by identifying the choice of $\zeta_{t}$ at any time $t\in\mathbb{N}$ as an action, and letting the state of the MFC MDP be given by the law of the time-variant system $(G(t),\mathbf{z}(t))$ resulting from the application of $\zeta_{t}$ when starting at some initial law.
@@ 1.3.3 Theoretical guarantees
Corollary 1.1.
Consider a finite set $\{\pi_{1},\ldots,\pi_{K}\}$ of MFC policies with MFC objective values $J^{G,\mathbf{z}}(\pi_{1}),\ldots,J^{G,\mathbf{z}}(\pi_{K})$. Let $\pi_{i}$ be the policy with maximal MFC objective value $J^{G,\mathbf{z}}(\pi_{i})\geq\max_{j}J^{G,\mathbf{z}}(\pi_{j})$. Then, there exists $n^{\prime}$ such that for all $n>n^{\prime}$, we also have optimality in the finite system
@@
The proofs are provided in the appendix and use the theoretical framework of Lacker et al. (2023) for general dynamical systems. In our experiments, we will also allow for randomized assignments per packet to further improve performance empirically, such that all packets arriving at a particular scheduler during the entirety of any epoch are independently randomly either allocated to the local queue or a random neighbour, i.e. formally we replace offloading choices $\mathcal{A}=\{0,1\}$ by probabilities for offloading each packet independently, $\mathcal{A}=[0,1]$.
The MFC MDP formulation and theoretical guarantees give us the opportunity to use MFC together with single-agent control such as RL in order to find good scalable solutions while circumventing hard exact analysis and improving over powerful techniques such as JSQ in large queueing systems. All that remains is to solve the limiting MDP.
1.4 Reinforcement Learning
Building upon the preceding MFC formulation, all that remains to find optimal load balancing in large sparse graphs, is to solve the MFC problem. Due to its complexity, the limiting MFC problem will be solved by considering it as a variant of an MDP, i.e. a standard formulation for single-agent centralized RL, which will also allow a model-free design methodology. More precisely, we will consider partially-observed MDP (POMDP) variant of the problem, since at any time $t$, we cannot evaluate the potentially infinite system $\mathcal{L}(G(t),\mathbf{z}(t))$ exactly to obtain action $\zeta_{t}=\pi(\mathcal{L}(G(t),\mathbf{z}(t)))$. Instead, we will use the empirical distribution $\mu^{G_{n},\mathbf{z}_{n}}(t)$ as an observation that is only correlated with the state of the entire system, but of significantly lower dimensionality ($|\mathcal{Z}|$-dimensional vector, instead of $\mathcal{Z}^{|G|}$ plus additional topological information). This also means that we need not consider the limiting system of potentially infinite size, or include the information of root nodes when considering network topologies in the following, which is intuitive as there is no notion of global root in local queueing systems.

### 22 Mar 2024 v2

@@ 2.2.4. Dynamical system model
The state of the agent is the state of its associated queue, and is affected by the actions of itself and neighbouring agents. By Poisson thinning, at every epoch $t\in\mathbb{N}$, given the current global state $\mathbf{z}(t)$ and action $\mathbf{a}(t)$, the next local state $z_{i}(t+1)$ of agent $i$ can be calculated independently only depending on the current state $\mathbf{z}_{i}(t)$ and actions $\mathbf{a}_{N_{i}}(t)\equiv\{a_{j}(t)\}_{j\in N_{i}}$, i.e. $\mathbb{P}(\mathbf{z}(t+1)\mid\mathbf{z}(t),\mathbf{a}(t))=\prod_{i=1}^{N}\mathbb{P}(z_{i}(t+1)\mid z_{i}(t),a_{N_{i}}(t))$, where each $\mathbb{P}(z_{i}(t+1)\mid z_{i}(t),a_{N_{i}}(t))$ can be computed by the Kolmogorov equation for continuous-time Markov chains, given that the rate of an arrival at a queue is given by the sum of arrival rates assigned by the agents, $\lambda_{t}^{i}=\lambda(t)(1-a_{i}+\sum_{j\in N_{i}}\frac{a_{j}}{|N_{j}|})$ such that the total arrival rate is $\sum_{i}\lambda_{t}^{i}=N\lambda(t)$, and the package departure rate of each queue is fixed to the serving rate $\alpha>0$.
@@
Each agent is associated with a local stage reward $r_{i}(z_{i},a_{i})$ and according global stage reward $r(\mathbf{z},\mathbf{a})=\frac{1}{N}\sum_{i=1}^{N}r_{i}(z_{i},a_{i})$. The reward is given in terms of a penalty for job drops due to each action $a_{i}$. The objective is then to find the localized policy tuple $\theta$ such that the global reward is maximized, starting from initial state distribution $z_{i}(0)\sim\mu_{0}$.
2.3. Sparse Graph Mean Field System
In this section, we will consider limits of large graphs and the behaviour in such systems, by using mean field theory from (Lacker et al., 2023) which focuses on an abstract theoretical framework, that is (i) unrelated to particular applications and (ii) does not explicitly consider control, which we integrate into the states of their discrete-time model.
@@ 2.3. Sparse Graph Mean Field System
In order to obtain a limiting mean field system, we assume a shared policy $\zeta_{i}=\zeta$ for all agents $i$. This assumption is natural, as it often gives state-of-the-art performance in various RL benchmarks (Yu et al., 2022; Christianos et al., 2020; Papoudakis et al., 2021) and also allows immediate application to arbitrary sizes of the finite system, as all schedulers can use the same policy. Furthermore, it will allow to scale up the system at any time by simply adding more schedulers and queues, without retraining a policy.
@@
In contrast to typical mean field games (Lasry and Lions, 2007) and mean field control (Bensoussan et al., 2013), we cannot reduce the entire system exactly to a single representative agent and its probability law. This is because in a sparse setting, the neighbourhood state distribution of any agent remains an empirical distribution of finitely many neighbours and hence cannot follow a law of large numbers into a deterministic mean field distribution. Therefore, the neighbourhood and its state distribution remain stochastic and it is necessary to model the probability law of entire neighbourhoods. The modelling of such graphical neighbourhoods is formalized in the following.
2.3.1. Topological structure
@@
To make sense of the limiting topology of our system formally, we introduce technical details, letting finite systems be given by sequences of (potentially random) finite graphs and initial states $(G_{n},\mathbf{z}_{n})$ converging in probability in the local weak sense to some limiting random graph $(G,\mathbf{z})$. Here, we assume that graphs are of bounded degree, i.e. there exists a finite degree $d$ such that all nodes have at most $d$ neighbours. In other words, we define a sequence of systems of increasing size according to a certain topology, which formalizes the scalable architectural choice of a network structure, such as a ring topology or torus. In the following paragraph, we give the formal definition, which may be skimmed by the reader.
@@
For convergence in the local weak sense, define first the space of marked rooted graphs $\mathcal{G}_{*}$, the elements of which essentially constitute a snapshot of the entire system at any point in time. Such a marked rooted graph consists of a tuple $(G,\varnothing,\mathbf{z})\in\mathcal{G}_{*}$, where $G=(\mathcal{N},\mathcal{E})$ is a graph, $\varnothing\in\mathcal{N}$ is a particular node of $G$ (the so-called root node), and $z\in\mathcal{Z}^{\mathcal{N}}$ defines states (”marks”) for each node in $G$, i.e. the current queue filling of queues associated to any agent (node). Denote by $B_{k}(G,\varnothing,\mathbf{z})$ the marked rooted subgraph of vertices in the $k$-hop neighbourhood of the root node $\varnothing$. The space $\mathcal{G}_{*}$ is metrized such that sequences $(G_{n},\varnothing_{n},\mathbf{z}_{n})\to(G,\varnothing,\mathbf{z})\in\mathcal{G}_{*}$ whenever for any $k\in\mathbb{N}$, there exists $n^{\prime}$ such that for all $n>n^{\prime}$ there exists a mark-preserving isomorphism $\phi\colon B_{k}(G_{n},\varnothing_{n},\mathbf{z}_{n})\to B_{k}(G,\varnothing,\mathbf{z})$, i.e. with $z_{ni}=z_{(\phi(i))}$ for all nodes $i\in G_{n}$ (local convergence). We will abbreviate elements $(G_{n},\varnothing_{n},\mathbf{z}_{n}),(G,\varnothing,\mathbf{z})\in\mathcal{G}_{*}$ as $(G_{n},\mathbf{z}_{n}),(G,\mathbf{z})$, and their node sets as $G_{n},G$ whenever it is clear from the context. Then, finally, convergence in the local weak sense is formally defined by $\lim_{n\to\infty}\frac{1}{|G_{n}|}\sum_{i\in G_{n}}f(C_{i}(G_{n}))=\mathbb{E}\left[f(G)\right]$ in probability for every continuous and bounded $f\colon\mathcal{G}_{*}\to\mathbb{R}$, where $C_{i}(G_{n})$ denotes the connected component of $i$ in $G_{n}$.
@@
In other words, wherever we randomly look in the graph, there will be little difference between the distribution of the random local system state (including its topology), and of the limiting $G$. This holds true, e.g. if we initialize all queues as empty or i.i.d., and consider certain types of topologies such as grids, see Section 3.1. More details also in (Lacker et al., 2023).
2.3.2. System model
@@
As discussed in the prequel, consider sequences of possibly random rooted marked graphs (i.e. finite graphs and initial states) $(G_{n},\mathbf{z}_{n})\to(G,\mathbf{z})$ with $|G_{n}|$ agents $i=\{1,\ldots,|G_{n}|\}$, converging in the local weak sense to the potentially infinite-sized system $(G,\mathbf{z})$. For a moment, consider a decentralized, stochastic control policy $\zeta\colon\mathcal{Z}\times\mathcal{T}\to[0,1]$ such that an agent $i\in G$ chooses to offload to a random neighbour (action $a_{i}=1$) with probability $\zeta_{t}(z_{i}(t))$, depending on its own queue state $z_{i}(t)$ only. We can then consider the probability law $\mathcal{L}(G,\mathbf{z})$ of the limiting system $(G,\mathbf{z})$ as the state of a single-agent MDP, similar to the MFC MDP formalism in standard MFC (Pham and Wei, 2018). This law is in essence the probability that the graph around any randomly chosen queue is in a certain state. At least formally, we do so by identifying the choice of $\zeta_{t}$ at any time $t\in\mathbb{N}$ as an action, and letting the state of the MFC MDP be given by the law of the time-variant system $(G(t),\mathbf{z}(t))$ resulting from the application of $\zeta_{t}$ when starting at some initial law.
@@ 2.3.3. Optimality guarantees
Corollary 2.2 ().
Consider a finite set $\{\pi_{1},\ldots,\pi_{K}\}$ of MFC policies with differing MFC objective values $J^{G,\mathbf{z}}(\pi_{1})$ to $J^{G,\mathbf{z}}(\pi_{K})$. Let $\pi_{i}$ be the policy with maximal MFC objective value $J^{G,\mathbf{z}}(\pi_{i})\geq\max_{j}J^{G,\mathbf{z}}(\pi_{j})$. Then, there exists $n^{\prime}$ such that for all $n>n^{\prime}$, we also have optimality in the finite system
@@
The proofs are provided in Appendix and use the theoretical framework of (Lacker et al., 2023) for general dynamical systems. In our experiments, we will also allow for randomized assignments per packet to further improve performance empirically, such that all packets arriving at a particular scheduler during the entirety of any epoch are independently randomly either allocated to the local queue or a random neighbour, i.e. formally we replace offloading choices $\mathcal{A}=\{0,1\}$ by probabilities for offloading each packet independently, $\mathcal{A}=[0,1]$.
The MFC MDP formulation and theoretical guarantees give us the opportunity to use MFC together with single-agent control such as RL in order to find good scalable solutions while circumventing hard exact analysis and improving over powerful techniques such as JSQ in large queueing systems. All that remains is to solve the limiting MDP, as MFC has formally converted MARL into single-agent RL.
2.4. Reinforcement Learning
Building upon the preceding MFC formulation, all that remains to find optimal load balancing in large sparse graphs, is to solve the MFC problem. Due to its complexity, the limiting MFC problem will be solved by considering it as a variant of an MDP, i.e. a standard formulation for single-agent centralized RL, which will also allow a model-free design methodology. Nonetheless, the training is still performed on the preceding finite graph multi-agent system, as we cannot evaluate the typically infinite limiting system $\mathcal{L}(G(t),\mathbf{z}(t))$.
More precisely, we will consider a partially-observed MDP (POMDP) variant of the problem, since at any time $t$, we cannot evaluate the potentially infinite system $\mathcal{L}(G(t),\mathbf{z}(t))$ exactly to obtain action $\zeta_{t}=\pi(\mathcal{L}(G(t),\mathbf{z}(t)))$. Instead, we will use the empirical distribution $\mu^{G_{n},\mathbf{z}_{n}}(t)$ as an observation that is only correlated with the state of the entire system, but of significantly lower dimensionality ($|\mathcal{Z}|$-dimensional vector, instead of $\mathcal{Z}^{|G|}$ plus additional topological information). This also means that we need not consider the limiting system of potentially infinite size, or include the information of root nodes when considering network topologies in the following, which is intuitive as there is no notion of global root in local queueing systems.